OpenStudy (anonymous):

Write the equation of the circle for the given center and radius. (-3,6) and radius of 6.

OpenStudy (anonymous):

(x+3)^2+(y-6)^2=36

OpenStudy (anonymous):

That's it? Then what's (x-h)^2+(y-k)^2=r^2 for?

OpenStudy (anonymous):

OH DUH nevermind

OpenStudy (anonymous):

Determine the coordinates of the center and the measure of the radius for the given equation. (x-4)^2+(y+1)^2=121 I got the center to be (4,-1) and the radius to be 11, is that right?